Suppose the number of vehicles crossing a bridge each day from 8am to 9am is meant to have a mean value 1500. The engineer counts the number of vehicles crossing the bridge for 25 days. The sample mean from the 25 observations is 1585, and the sample standard deviation is 162. Conduct a hypothesis test assuming an unknown standard deviation, to determine if the true mean is 1500 cars per day or not. Use α=0.05.
Answer #1

H0 : mu =1500
Ha: mu not = 1500

test statistics:

t = (xbar -mu)/(s/sqrt(n))
= ( 1585 - 1500)/(162/sqrt(25))
= 2.623


p value = 2*(1- P(t < 2.623)) with df =24

P value = 0.0149


As p value < 0.05 Rejct H0